A manatee travels 5 miles in 1 1/4 hr. How far will he travel in 1 hour?
Found 2 solutions by Alan3354, TimothyLamb:
Answer by Alan3354(69443)

A manatee travels 5 miles in 1 1/4 hr. How far will he travel in 1 hour?
--------------
It's 0.8 of the time --> 0.8 of the distance.

Answer by TimothyLamb(4379)

s = d / t
s = 5 / 1.25 = 4 mph
---
d = s * t
d = 4 * 1 = 4
Answer:
4 miles
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php